The Pappas family has two cats, Zeus and Athena. Together they weigh twelve pounds. Zeus weighs 4 pounds less than Athena. How much does Athena weigh?
Write an algebraic equation that represents the situation.
Use a for your variable.
Athena weighs pounds.

Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52778)

.

    a + (a-4) = 12

    2a - 4 = 12

    2a = 12 + 4

    2a = 16

     a = 16/2 = 8.


ANSWER.  Athena weights 8 pounds.

Solved and presented in a form of a mantra to repeat it
many times again and again for any other similar problem.


//////////////////////////


It always touches me when I see a request "solve the problem by introducing variables in such and such a way."
Then I think where these people fell from.


Real teachers usually avoid to give instructions in Math at such micro-managing level.
Micro-managing in Math is the way to create a generations of students who will hate Math.



Answer by josgarithmetic(39617)

Zeus                  x-4

Athena                 x

COMBINED              12 

With that, should become both obvious and simple.